Find the new equation of the curve (x-2)^2= y(y-1)^2 by the transforming to parallel axes through the point (2,1)
1
Expert's answer
2020-11-18T18:31:06-0500

"(x-2)^2 = y(y-1)^2"

Let the point on the curve with respect to old axis is (x,y)

Let the same point on the curve with respect to new axis be (x ′ ,y ′ )

The new co-ordinates are shifted by (2,1) Thus x=x ′ +2 and y=y ′ +1

Replacing the value of x and y in "(x-2)^2 = y(y-1)^2" ,we get

"(x'+2-2)^2 = (y'+1)(y'+1-1)^2\\
(x')^2 = (y'+1)(y')^2"

answer: "(x')^2 = (y'+1)(y')^2"